# Trade-off of Security, Latency, and Throughput of the Nakamoto Consensus

### v3

@@ 5.3. A Sketch of the Statistical Analysis
For every integer $n\geq 0$, the number of balanced heights up to the $k$-th candidacy, denoted as $X^{(k)}$, satisfies
(35) $P(X^{(k)}\leq n)\geq F^{(k)}(n)$
where $F^{(k)}$ is defined in (14). As a special case, the total number of balanced heights defined in (28) satisfies
(36) $P(X\leq n)\geq F(n)$
where
@@
Theorem 2 is proved in Appendix B. The key insight here is that the adversary can only take advantage of a bounded number of H-blocks to balance each other even in the long run. In fact, with probability 1, one last balancing opportunity will be missed and the adversary is then behind future pacers permanently. The final balancing candidate serves the role of the “Nakamoto block” in (dembo2020everything, ) and is a stronger notion than the “convergence opportunity” in e.g., (pass2017rethinking, ). This observation underlies the safety guarantee of Theorem 2.
@@
(37) $Y^{(k)}=L_{s}-k+\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}.$
While $Y^{(k)}$ and $X_{s,\tau}$ are dependent, we can upper bound the probability of the event (29), i.e., $X_{s,\tau}+Y^{(k)}\geq 0$, using a union bound:
@@
It suffices to upper bound the marginal complementary cdf’s of $X_{s,\tau}$ and $Y^{(k)}$. Since all $X_{s,\tau}$ balanced heights must be within the first $k$ heights above $h_{b}$, we have
(39) $P(X_{s,\tau}>n-1)\leq P(X^{(k)}>n-1)$
@@
$P(Y^{(k)}$ $+X_{s,\tau}\geq 0)$
(40) $\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X_{s,\tau}}(n-1)+\bar{F}_{Y^{(k)}}(-n)\}$
(41) $\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X^{(k)}}(n-1)+\bar{F}_{Y^{(k)}}(-n)\}.$
@@
(42) $J_{s,t}\geq P_{s+\Delta,t},\quad\forall t>s.$
@@
(43) $Y^{(k)}$ $\leq L_{s}-k+\sup_{d\geq\tau}\{A_{s,d}-P_{\tau+\Delta,d-\Delta}\}$
(44) $=L_{s}-k+\sup_{c\geq 0}\{A_{s,\tau+2\Delta+c}-P_{\tau+\Delta,\tau+\Delta+c}\}.$
@@ 5.4. Operational Implications
A brief discussion on the operational implications of the theorems presented in Sec. 2 is warranted. First, the requirement for “a sufficiently large $s$” in the theorems serves the purpose of ensuring that the maximum random pre-mining lead is in a steady state. This assumption is easily applicable in most practical scenarios.
@@
More importantly, we define $s$ as “given” in the sense that it is not influenced by mining processes. Specifically, it is not selected by the adversary. If the adversary were to choose $s$, they could make a target transaction public once they have established a substantial lead (this might take an exceedingly long time, but it would eventually occur with a probability of 1). By doing so, a safety violation would be guaranteed. In this paper, we eliminate such adaptive attacks by assuming that the transaction becomes public within a fixed time window that is beyond the adversary’s control.
Operationally, we have the flexibility to select an arbitrary number of intervals $(s,s+2\Delta]$. If an honest node first encounters a transaction during $(s,s+\Delta]$ (who will also propagate this transaction), then the node can be sure that the transaction becomes public during the interval $(s,s+2\Delta]$. The node can be confident about the probabilistic safety guarantees for the transaction due to the theorems given in Sec. 4.
@@
Following the preceding outline of the statistical analysis, we prove the main theorems in detail in the appendices. In Sec. 6, we present the numerical results derived from these theorems. These results demonstrate the practicality of our latency-security trade-off bounds and their utility as guidance of parameter selection for Nakamoto-style protocols.
6. Numerical Result
@@
In this section, the system parameters are chosen to represent two different settings in most cases. The first setting is modeled after Bitcoin, where the total mining rate $h+a=1/600$ blocks/second. Bitcoin blocks’ propagation delays vary and have been decreasing over the years. As of October 2023, the 90th percentile of block propagation delay is about 4 seconds on average (bitcoinmoniter, ). We use $\Delta=10$ seconds as an upper bound. The second setting is modeled after Ethereum Classic (ETC), where $h+a=1/13$ blocks/second and $\Delta=2$ seconds. We also include results for Litecoin and Dogecoin parameters in one table.
6.1. Security Latency Trade-off
Figure 3. Bitcoin’s safety versus confirmation time.
Figure 4. ETC’s safety versus confirmation time.
@@ 6.1. Security Latency Trade-off
In Fig. 3, we first show the achievable and unachievable probability of safety violation as a function of the confirmation time (in hours) under the Bitcoin setting. For three different adversarial mining fractions ($\beta=1/10$, $1/7$, and $1/4$), we plot the upper and lower bounds calculated using Theorems 8 and 9, respectively. Also included for comparison is an achievable trade-off from (li2021close, , Eqn. (5)).
In Fig. 4, we plot the latency-security trade-off of ETC for two different adversarial mining fractions ($\beta=1/7$ and $1/4$). The gap between the upper and lower bounds is smaller than the gap derived in (li2021close, ).
@@
In Fig. 5, we show Bitcoin’s safety bounds as a function of confirmation depth $k$ calculated using Theorems 25, and 6, respectively. Also included are the achievable and unachievable bounds from (guo2022bitcoin, , Eqns. (7) and (8)) for comparison. The new upper bound (5) of Theorem 5 yields substantially tighter result. Similarly, ETC’s safety bounds as a function of confirmation depth $k$ is plotted in Fig. 6. For the ETC parameters, the achievable bounds in this paper significantly outperform the achievable bound of (guo2022bitcoin, , Eqn. (7)).
Figure 5. Bitcoin’s safety versus confirmation depth.
Figure 6. ETH’s safety versus confirmation depth.
@@
$\beta=10\%$ $7$ $7$ $7$ $10$ $21$ $52$
$\beta=20\%$ $14$ $14$ $14$ $26$ $35$ $117$
$\beta=30\%$ $34$ $35$ $35$ $63$ $91$ $633$
$\beta=40\%$ $150$ $154$ $161$ $277$ $552$infty
Table 2. The confirmation depth needed to guarantee $99.9\%$ safety for various block delays and adversarial fractions.
For a fixed $99.9\%$ safety guarantee, using the Bitcoin parameters, we compute the confirmation depth needed according to Theorem 5 for various block propagation delays and adversarial mining fractions. Table 2 shows that the confirmation depth rises slowly with the delay bound when the adversarial fraction is relatively low; it rises more quickly when the adversarial fraction is relatively high.
@@
Table 3 shows the probability of having more than a given number of balanced heights with 25% adversarial mining. For example, the probability of having more than four balanced heights is less than $10^{-9}$ for Bitcoin. Hence the bound in Lemma 3 is fairly tight for our purposes.
$\lambda^{-1}$ (s) $\Delta$ (s) $10^{-3}$ $10^{-6}$ $10^{-9}$
Bitcoin (guo2022bitcoin, ) $600$ 10 $\leq 25$ $\leq 51$ $\leq 78$
Bitcoin new $600$ 10 $\leq 23$ $\leq 47$ $\leq 71$
Litecoin (guo2022bitcoin, ) $150$ 10 $\leq 36$ $\leq 74$ $\leq 110$
Litecoin new $150$ 10 $\leq 32$ $\leq 63$ $\leq 99$
Dogecoin (guo2022bitcoin, ) $60$ 10 $\leq 85$ $\leq 171$ $\leq 258$
Dogecoin new $60$ 10 $\leq 52$ $\leq 127$ $\leq 211$
ETC (guo2022bitcoin, ) $13$ 2 $\leq 73$ $\leq 151$ $\leq 230$
ETC new $13$ 2 $\leq 47$ $\leq 114$ $\leq 193$
@@
Table 4. Bounds of confirmation depth (latency) for different security levels with $25\%$ adversarial mining power.
@@
Table 4 presents the depth needed to guarantee different security levels (probability of target transaction’s safety being violated) including $10^{-3}$, $10^{-6}$, $10^{-9}$, for Bitcoin, Litecoin, Dogecoin, and ETC, respectively. The table shows that the new upper bound due to Theorem 5 is often substantially smaller than the previous upper bound from (guo2022bitcoin, ).
6.2. Confirmation versus Fault Tolerance
@@
In this subsection, we plot the confirmation depth or time that guarantees a desired safety level as a function of the fraction of adversarial mining, while holding other parameters fixed. This also makes it easier to see the maximum fault tolerance different bounding methods can apply to. We assume the ETC parameters.
Refer to captionRefer to caption
@@
Figure 7. Confirmation latency that guarantees 99.9% safety for ETC assuming different fractions of adversarial mining. Top: Confirmation in depth. Bottom: Confirmation in time (minutes).
@@
Fig. 7 shows new and previous bounds on the confirmation depth/time that guarantees 99.9% safety as a function of the adversarial fraction of mining power. The top graph uses confirmation by depth (Theorems 25, and 6), whereas the bottom graph uses confirmation by time (Theorems 8 and 9). We include two points reported in (gazi2022practical, ), where the numerical evaluation becomes impractical for much higher adversarial mining rates. Also included in the graph are achievable infinite-latency fault tolerances established in (garay2015bitcoin, )(pass2017analysis, )(li2021close, ) and the ultimate tolerance established in (dembo2020everything, )(gazi2020tight, ), and also in this paper. The new upper and lower bounds are significantly closer than all previous bounds. Even our new closed-form upper bound is quite competitive.
6.3. Throughput-Latency Trade-Off
@@
In this subsection, we evaluate the best throughput we can guarantee for optimized parameters including block size, mining rate, and confirmation depth. We fix the safety guarantee to be $99.9\%$. We guarantee up to $\beta=40\%$ adversarial mining, which implies that $\lambda\Delta<5/6$ due to (19). We assume the block propagation delay and the block size satisfy the affine relationship (21). Using empirically determined delay bounds of 10 seconds for 1.6 MB Bitcoin blocks and 2 seconds for 200 KB Bitcoin Cash blocks, [^3] The 90th percentile of Bitcoin Cash propagation time is 1 second (fechner2022calibrating, ). we infer that $r=178.6$ KB/second and $\nu=0.9$ (seconds).
@@
Figure 8. Achievable and unachievable throughput-latency trade-off.
@@
6.4. Guidance to Selecting Parameters
@@
The theorems in this paper provide the most accurate guidance to date to selecting parameters for desired system-level performance. For example, if the block size is increased to $2$ MB, the propagation upper bound will be around $\Delta=12.5$ seconds according to (21). If the total mining rate is increased to $\lambda=1/30$ block/seconds, the throughput increases to over $40$ KB/s according to (23a). At $\lambda\Delta=5/12$, the ultimate fault tolerance is $44.44\%$ adversarial mining; in this case, at $30$% adversarial mining, the latency for achieving $10^{-6}$ safety is less than $3.5$ hours, while original Bitcoin takes more than $18$ hours. Thus the throughput, latency, and safety are simultaneously improved for a minor sacrifice in the fault tolerance. In fact, $44.44\%$ adversarial mining is practically not tolerable using the original Bitcoin parameters due to the excessively long delays of thousands of block confirmations.
7. Conclusion
In this paper, we have developed a concise mathematical model for blockchain systems based on the Nakamoto consensus. We have obtained a new set of closed-form latency-security bounds, which are the first such bounds to apply to all parameter values. These bounds are not only easily computed but also demonstrate superior numerical performance across a broad spectrum of parameter configurations. Our analysis not only provides an intuitive proof of the ultimate fault tolerance but also offers new insights for the practical application the Nakamoto consensus.
Existing safety guarantees for longest-chain-fork-choice protocols have primarily been established through “block counting” methods, which calculates, the probability that one species of blocks ismined more than another species during some periods longer than the confirmation latency. Such analyses essentially empower the adversary to use non-causal knowledge of future blocks (up until confirmation) and advance them to build a winning chain. In contrast, the production of a balanced height depends on blocks mined within an opportunity window of $\Delta$ seconds; future blocks cannot be used to create balancing candidates or balanced heights. This requirement eliminates certain non-causal attack strategies, resulting in improvements and the precise determination of fault tolerance.

### v5

@@ V-C A Sketch of the Statistical Analysis
With respect to a given target block, the number of balanced heights up to the $k$-th candidacy, denoted as $X^{(k)}$, satisfies
$P(X^{(k)}\leq n)\geq F^{(k)}(n)$ (46)
for every integer $n\geq 0$, where $F^{(k)}$ is defined in (15). As a special case, the total number of balanced heights defined in (29) satisfies
$P(X\leq n)\geq F(n).$ (47)
@@
Theorem 11 is proved in Appendix B. The key insight here is that the adversary can only take advantage of a bounded number of H-blocks to balance each other even in the long run. In fact, with probability 1, one last balancing opportunity will be missed and the adversary is then behind future pacers permanently. The final balancing candidate serves the role of the “Nakamoto block” in [3] and is a stronger notion than the “convergence opportunity” in e.g., [14]. This observation underlies the safety guarantee of Theorem 2.
@@
$Y^{(k)}=L_{s}-k+\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}.$ (48)
While $Y^{(k)}$ and $X_{t_{f}+\Delta}^{g}$ are dependent, we can upper bound the probability of the event (30), i.e., $X_{t_{f}+\Delta}^{g}+Y^{(k)}\geq 0$, using a union bound:
@@
It suffices to upper bound the marginal complementary cdf’s of $X_{t_{f}+\Delta}^{g}$ and $Y^{(k)}$. Since all $X_{t_{f}+\Delta}^{g}$ balanced heights must be within the first $k$ heights above $h_{b}$, we have
$P(X_{t_{f}+\Delta}^{g}>n-1)\leq P(X^{(k)}>n-1)$ (50)
@@
$P(Y^{(k)}$ $+X_{t_{f}+\Delta}^{g}\geq 0)$
$\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X_{t_{f}+\Delta}^{g}}(n-1)+\bar{F}_{Y^{(k)}}(-n)\}$ (51)
$\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X^{(k)}}(n-1)+\bar{F}_{Y^{(k)}}(-n)\}.$ (52)
@@
$J_{s,t}\geq P_{s+\Delta,t},\quad\forall t>s.$ (53)
@@
$Y^{(k)}$ $\leq L_{s}-k+\sup_{d\geq\tau}\{A_{s,d}-P_{\tau+\Delta,d-\Delta}\}$ (54)
$=L_{s}-k+\sup_{c\geq 0}\{A_{s,\tau+2\Delta+c}-P_{\tau+\Delta,\tau+\Delta+c}\}.$ (55)
@@ V-D Operational Implications
A brief discussion on the operational implications of theorems presented in Sec. 2 is warranted. First, the requirement for “a sufficiently large $s$” in the theorems serves the purpose of ensuring that the maximum random pre-mining lead is in a steady state. This assumption is easily applicable in most practical scenarios.
@@
More importantly, we define $s$ as “given” in the sense that it is not influenced by mining processes. Specifically, it is not selected by the adversary. If the adversary were to choose $s$, it could choose a target block one height above the highest public height once they have established a substantial lead (this might take an exceedingly long time, but it would eventually occur with a probability of 1). By doing so, a safety violation would be guaranteed. In this paper, we eliminate such adaptive attacks by assuming that the target block needs to extend a credible chain that is beyond the adversary’s control.
Operationally, we have the flexibility to select an arbitrary time $s$. If an honest node sees the target block mined on top of a highest block in its view at time $s$, the node can be sure that the block is mined on $s$-credible chain, and hence satisfies the conditions in the theorems given in Sec. IV.
@@
Following the preceding outline of the statistical analysis, we prove the main theorems in detail in the appendices. In Sec. VI, we present the numerical results derived from these theorems. These results demonstrate the practicality of our latency-security trade-off bounds and their utility as guidance of parameter selection for Nakamoto-style protocols.
VI Numerical Result
@@
In this section, the system parameters are chosen to represent two different settings in most cases. The first setting is modeled after Bitcoin, where the total mining rate $\lambda=1/600$ blocks/second. Bitcoin blocks’ propagation delays vary and have been decreasing over the years. As of November 2024, the 90th percentile of block propagation delay is about 4 seconds on average [24]. We use $\Delta=10$ seconds as an upper bound. The second setting is modeled after Ethereum Classic (ETC), where $\lambda=1/13$ blocks/second and $\Delta=2$ seconds. We also include results for Litecoin and Dogecoin parameters in one table.
VI-A Security Latency Trade-off
Figure 4: Bitcoin’s safety versus confirmation time.
Figure 5: ETC’s safety versus confirmation time.
@@ VI-A Security Latency Trade-off
In Fig. 4, we first show the achievable and unachievable probability of safety violation as a function of the confirmation time (in hours) for the Bitcoin configuration. For three different adversarial mining fractions ($\beta=1/10$, $1/7$, and $1/4$), we plot the upper and lower bounds calculated using Theorems 8 and 9, respectively. Also included for comparison is an achievable trade-off from [6, Eqn. (5)]. In Fig. 5, we plot the latency-security trade-off of ETC for two different adversarial mining fractions ($\beta=1/7$ and $1/4$). In all cases, the new bounds approximately halve the latency gaps derived in [6].
@@
In Fig. 6, we show Bitcoin’s safety bounds as a function of confirmation depth $k$ calculated using Theorems 25, and 6, respectively. Also included are the achievable and unachievable bounds from [8]. The new upper bound (18) of Theorem 5 yields tighter results for small confirmation depths. Similarly, ETC’s safety bounds as a function of confirmation depth $k$ is plotted in Fig. 7. For the ETC parameters, the achievable bounds in this paper outperform the achievable bound of [8].
Figure 6: Bitcoin’s safety versus confirmation depth.
Figure 7: ETH’s safety versus confirmation depth.
@@
$\beta=10\%$ $7$ $7$ $8$ $15$ $16$ $37$
$\beta=20\%$ $14$ $16$ $16$ $19$ $28$ $65$
$\beta=30\%$ $34$ $37$ $39$ $46$ $65$ $633$
$\beta=40\%$ $150$ $153$ $156$ $277$ $552$infty
TABLE III: The confirmation depth needed to guarantee $99.9\%$ safety for various block delays and adversarial fractions.
For a fixed $99.9\%$ safety guarantee, using the Bitcoin parameters, we compute the confirmation depth needed according to Theorem 5 for various block propagation delays and adversarial mining fractions. Table III shows that the confirmation depth rises slowly with the delay bound when the adversarial fraction is relatively low; it rises more quickly when the adversarial fraction is relatively high.
@@
Table IV shows the probability of having more than a given number of balanced heights with 25% adversarial mining. For example, the probability of having more than four balanced heights is less than $10^{-9}$ for Bitcoin. Hence the bound in Lemma 12 is fairly tight for our purposes. We note that for small confirmation depths, the probability of mining more adversarial blocks than pacers dominates the probability of balancing any height. As the confirmation depth increases, the latter probability begins to make a noticeable impact, hence the bump on the curves in Fig. 6.
$\lambda^{-1}$ (s) $\Delta$ (s) $10^{-3}$
Bitcoin [7] $600$ 10 $\leq 25$
Bitcoin new $600$ 10 $\leq 24$
Litecoin [7] $150$ 10 $\leq 36$
Litecoin new $150$ 10 $\leq 32$
Dogecoin [7] $60$ 10 $\leq 85$
Dogecoin new $60$ 10 $\leq 52$
ETC [7] $13$ 2 $\leq 73$
ETC new $13$ 2 $\leq 47$
@@
TABLE V: Bounds of confirmation depth (latency) to guarantee $99.9\%$ safety with $25\%$ adversarial mining power.
@@
Table V presents the depth needed to guarantee a $99.9\%$ for Bitcoin, Litecoin, Dogecoin, and ETC, respectively. The table shows that the new upper bound due to Theorem 5 is often substantially smaller than the previous upper bound from [7].
VI-B Confirmation versus Fault Tolerance
@@
In this subsection, we plot the confirmation depth or time that guarantees a desired safety level as a function of the fraction of adversarial mining, while holding other parameters fixed. This also makes it easier to see the maximum fault tolerance different bounding methods can be applied to. We base our analysis on the parameters of Ethereum Classic (ETC).
@@
Figure 8: Confirmation latency that guarantees 99.9% safety for ETC assuming different fractions of adversarial mining. Top: Confirmation in depth. Bottom: Confirmation in time (minutes).
@@
Fig. 8 shows new and previous bounds on the confirmation depth/time that guarantees 99.9% safety as a function of the adversarial fraction of mining power. The top graph uses confirmation by depth (Theorems 25, and 6), whereas the bottom graph uses confirmation by time (Theorems 8 and 9). We include two points reported in [5], where the numerical evaluation becomes impractical for much higher adversarial mining rates. Also included in the graph are achievable infinite-latency fault tolerances established in [6][7][8][11][13], and the ultimate tolerance established in [3][4], and also in this paper. The new upper and lower bounds are significantly closer than all previous bounds. Even our new closed-form upper bound is highly competitive.
VI-C Throughput-Latency Trade-Off
@@
In this subsection, we evaluate the best throughput we can guarantee for optimized parameters including block size, mining rate, and confirmation depth. We fix the safety guarantee to be $99.9\%$. We guarantee up to $\beta=40\%$ adversarial mining, which implies that $\lambda\Delta<5/6$ due to (20). We assume the block propagation delay and the block size satisfy the affine relationship (22). Using empirically determined delay bounds of 10 seconds for 1.6 MB Bitcoin blocks and 2 seconds for 200 KB Bitcoin Cash blocks, [^3] The 90th percentile of Bitcoin Cash propagation time has been measured as 1 second [12]. we infer that $r=178.6$ KB/second and $\nu=0.9$ (seconds).
@@
Figure 9: Achievable and unachievable throughput-latency trade-off and 5 existing currencies’ throughput-latency.
@@
VI-D Guidance for Selecting Parameters
@@
The theorems in this paper provide the most accurate guidance to date for selecting parameters for desired system-level performance. For example, if the block size is increased to $2$ MB, the propagation upper bound will be around $\Delta=12.5$ seconds according to (22). If the total mining rate is increased to $\lambda=1/30$ block/seconds, the throughput increases to over $40$ KB/s according to (24a). At $\lambda\Delta=5/12$, the ultimate fault tolerance is $44.44\%$ adversarial mining; in this case, at $30$% adversarial mining, the latency for achieving $10^{-6}$ safety is less than $3.5$ hours, while original Bitcoin takes more than $18$ hours. Thus the throughput, latency, and safety are simultaneously improved for a minor sacrifice in the fault tolerance. In fact, $44.44\%$ adversarial mining is practically not tolerable using the original Bitcoin parameters due to the excessively long delays of thousands of block confirmations.
VII Conclusion
In this paper, we have developed a concise mathematical model for blockchain systems based on the Nakamoto consensus. We have obtained a new set of closed-form latency-security bounds, which are the first such bounds to achieve the ultimate fault tolerance. These bounds are not only easily computed but also demonstrate superior numerical performance across a broad spectrum of parameter configurations. Our analysis not only provides an intuitive proof of the ultimate fault tolerance but also offers new insights for the practical application the Nakamoto consensus.
Existing safety guarantees for longest-chain fork-choice protocols have primarily been established through “block counting” methods, which calculates, the probability that one species of blocks ismined more than another species during some periods longer than the confirmation latency. Such analyses essentially empower the adversary to use non-causal knowledge of future blocks (up until confirmation) and advance them to build a winning chain. In contrast, the production of a balanced height depends on blocks mined within an opportunity window of $\Delta$ seconds; future blocks cannot be used to create balancing candidates or balanced heights. This requirement eliminates certain non-causal attack strategies, resulting in improvements and achieving the ultimate fault tolerance.